fran swims at a speed of 2.2 km/h in still water. the lazy river flows at a speed of 0.4 km/h. how long will it take fran to swim 0.9 km upstream?
Answer by nerdybill(7384)

fran swims at a speed of 2.2 km/h in still water. the lazy river flows at a speed of 0.4 km/h. how long will it take fran to swim 0.9 km upstream?
.
We will need to apply the "distance" formula:
d = rt
where
d is distance
r is rate or speed
t is time
.
Reordering the equation, we solve for 't':
t = d/r
.
Fran's rate (r) is:
2.2 - 0.4 = 1.8 km/h
.
Therefore:
.9/1.8 = 0.5 hours